### v1

# Enhanced superconductivity in LaRu3Si2 by chemical pressure tuning of kagome flat bands

Abstract
In kagome metals, flat bands induced by frustrated hopping serve as a platform for strong electronic correlation, while often exhibiting substantial dispersion along the $k_{z}$ direction due to interlayer coupling. Here, we investigate the superconductivity of kagome metal LaRu3(Si1-xGex)2 by chemical pressure tuning while preserving the Ru-$4d$ states that constitute the kagome flat bands. We observe a sizable enhancement in the density of states up to $x=0.07$, as determined by the specific heat, with a concomitant increase in the superconducting transition temperature. As supported by our first-principles calculations, the lattice expansion along the $c$-axis by chemical pressure mitigates a detrimental effect of hybridization between kagome layers, thereby reinforcing the three-dimensional localization of flat-band electrons. Providing evidence for LaRu3Si2 as a flat-band hosting superconductor, we demonstrate a simple, yet effective, route to engineer flat bands by uniaxial lattice expansion.
Main Text
The kagome lattice, a two-dimensional network of corner-sharing triangles, features three sublattice degrees of freedom per unit cell. kagome metals serve as a platform for investigating the intricate interplay between electron-electron correlations and the topology of electronic band structures. Yin et al. (2022); Wang et al. (2023); Checkelsky et al. (2024). This study discusses superconductivity in the intermetallic LaRu3Si2, where the kagome lattice is formed by Ru atoms as depicted in Fig. 1(a). Considering first a single orbital model on an isolated kagome lattice in Fig. 1(b), one obtains three bands forming a flat band and Dirac cone, and with van-Hove singularities at the $M$-points in the Brillouin zone [Fig. 1(c)]. In particular, the dispersionless upper band originates from frustrated hopping between the three sites of the kagome lattice Bergman et al. (2008). Further introducing Kane-Mele type spin-orbit coupling Kane and Mele (2005), a gap opens between the Dirac cone and the flat band, imparting a Chern number onto the flat band in this two-dimensional model Guo and Franz (2009a); Bolens and Nagaosa (2019); Kang et al. (2020).
Real-world kagome metals retain the features of this simplified electronic structure if the coupling between consecutive kagome layers is weak, and mostly $3d$ electron systems have been considered as realizations of this scenario Lin et al. (2018); Kang et al. (2019, 2020); Li et al. (2021); Regnault et al. (2022); Ye et al. (2024). As an example, Fig. 1(d) presents the electronic band structure of LaRu3Si2, which hosts an hourglass-shaped dispersion originating from Ru-$4d$ orbitals along the $\Gamma$$K$$M$ path. Early experimental work and a recent theoretical study emphasize that LaRu3Si2 distinguishes itself as the only known kagome superconductor hosting flat bands near the Fermi level without magnetic order Mielke et al. (2021); Junze et al. (2025). It is further suggested that the flat bands may be a driver of superconductivity with the highest observed transition temperature $T_{\mathrm{c}}=6.7\,$K among the family of $RT_{3}M_{2}$ ($R$ = rare-earth, $T$ = transition metal and $M$ = main group element) Barz (1980); Ku et al. (1980); Chaudhary et al. (2023); Gui and Cava (2022); Rauchschwalbe et al. (1984); Mielke et al. (2021); Gong et al. (2022) and among Kagome metals as a material class Deng et al. (2025). Moreover, charge-density-wave orders, possibly accompanied by breaking of time-reversal symmetry, have been recently observed Plokhikh et al. (2024); Mielke et al. (2024), reminiscent of the kagome superconductors $A$V3Sb5 ($A=$K, Cs, and Rb) Wilson and Ortiz (2024); Ortiz et al. (2019).
Real kagome metals are three-dimensional, and their flat bands often exhibit strong dispersion along the $k_{z}$ direction of reciprocal space Kang et al. (2020); Huang et al. (2022); Ye et al. (2024). This behavior is also observed in LaRu3Si2 along the $\Gamma-A$ line in Fig. 1(d). Therefore, flat-band electrons are not fully localized in known bulk kagome metals. To achieve tailored flat bands, it has thus been proposed to either suppress interlayer coupling [$t_{\mathrm{out}}$ in Fig. 1 (b)] by isolating kagome layers Jovanovic and Schoop (2022), or to introduce 3D frustrated hopping Wakefield et al. (2023). For the latter approach, 3D flat bands have been predicted for the pyrochlore lattice Guo and Franz (2009b), a 3D analogue of the kagome lattice, and have recently been observed Wakefield et al. (2023); Huang et al. (2024). Despite its simplicity, the former approach has not yet been conclusively implemented to enhance electron correlation phenomena and superconductivity on the kagome lattice.
Here, we engineer the electronic structure of LaRu3Si2 by substituting Si with the larger volume element Ge. We observe a marked increase of the superconducting critical temperature $T_{\mathrm{c}}$. An enhancement of the total density of states (DOS) with Ge-doping underpins the observed evolution of $T_{\mathrm{c}}$, as confirmed by specific heat measurements. Our electronic structure calculations demonstrate that anisotropic lattice expansion along the $c$-axis isolates the kagome layers, weakens the interlayer hopping $t_{\mathrm{out}}$, and promotes the DOS at the Fermi level while keeping the intralayer hopping $t$ intact. As summarized in Fig. 1(e), this is in stark contrast to previous attempts to tune superconductivity in LaRu3Si2 by chemical substitution at the Ru site, which generally results in a sharp decrease of $T_{\mathrm{c}}$ Li et al. (2012, 2016); Chakrabortty et al. (2023).
Crystals were obtained by the melting technique SI . In Fig. 2(a), the LaRu3Si2 structure was confirmed by powder X-ray diffraction (XRD). The absence of the LaRu2Si2 phase is confirmed. Figure 2(c,d) shows the hexagonal lattice constants of LaRu3(Si1-xGex)2(LRSG) extracted from XRD and their evolution with $x$ based on the LeBail analysis in the $P6/mmm$ structure. This neglects minor effects of symmetry-lowering charge order in this family of compounds Plokhikh et al. (2024). We find a robust value of $a$ and a linear expansion of the out-of-plane lattice constant $c$, which in combination gives a linear volume expansion in Fig. 2(e). The latter is consistent with Vegard’s law, and in good agreement with the trend derived from structure relaxation calculations of LaRu3Si2 and LaRu3Ge2. Magnetization and specific heat measurements were performed in a commercial cryostat SI .
To confirm the bulk superconducting state in LRSG, we show the magnetic susceptibility $\chi$ in Fig. 3(a). Above the transition temperature $T_{\mathrm{c}}$, $\chi$ is small and positive, consistent with the weak paramagnetic Pauli susceptibility of LaRu3Si2 Kishimoto et al. (2002). Strong diamagnetism accompanying the superconducting transition appears as a sharp drop in the data around $6\sim 7\,$K.
To precisely determine $T_{\mathrm{c}}$, we measure the specific heat $C(T)$ in zero field and at 555\,5T, which is high enough to entirely suppress the superconductivity [Fig. 3(b)]. In the former traces, a sharp kink appears at $T_{\mathrm{c}}$, the second-order transition into the superconducting state. We model the normal state contribution $C_{\mathrm{N}}$ by electronic and phonon terms, viz. $C_{\mathrm{N}}(T)=\gamma T+\beta T^{3}+\eta T^{5}$ [lines in Fig. 3(b)]. The higher order phonon contribution $\eta T^{5}$ is necessary to describe the data  Li et al. (2011). After subtraction of $C_{\mathrm{N}}(T)$, we fit the superconducting specific heat $C(T)-C_{\mathrm{N}}(T)$ to the prediction of Bardeen-Cooper-Schrieffer (BCS) theory by equal-area construction in Fig. 3(c) Tinkham (2004), confirming an enhancement in $T_{\mathrm{c}}$ with $x$. Next, we discuss the evolution of $\gamma$ and $\beta$ from the fit of $C_{\mathrm{N}}(T)$ in Fig. 3(d). A clear increase of $\gamma$ and $\beta$ is observed with $x$. Given the formula $\gamma=\pi^{2}k_{\mathrm{B}}^{2}N(E_{\mathrm{F}})/3$ with $N(E_{\mathrm{F}})$ being the density of states at the Fermi energy $E_{\mathrm{F}}$, the enhancement of $\gamma$ corresponds to an increase of $N(E_{\mathrm{F}})$ [Fig. 3(e)]. The trend is reproduced by DFT, where $N(E_{\mathrm{F}})$ increases from $2.9\,\mathrm{states/eV/u.c/spin}$ to $3.3\,\mathrm{states/eV/u.c/spin}$ as seen in Fig. 3(f). The marked enhancement of $N(E_{\mathrm{F}})$ goes hand in hand with an increase of $T_{\mathrm{c}}$, reaching up to 7.17.17.1\,7.1K for $x=0.07$ [Fig. 3(e)].
The increase of $N(E_{\mathrm{F}})$ and $T_{\mathrm{c}}$ qualitatively follows the McMillan equation
$T_{\mathrm{c}}=\frac{\Theta_{\mathrm{D}}}{1.45}\exp\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right),$ (1)
where $\Theta_{\mathrm{D}}$ is the Debye temperature, $\lambda$ is the electron-phonon coupling constant, and $\mu^{*}$ is the Coulomb pseudopotential, which usually takes a value around $0.1$ for a metal Carbotte (1990). Assuming a relatively static pairing potential $V$, the enhancement of $N(E_{\mathrm{F}})$ directly corresponds to that of $\lambda$, as given by the relation $N(E_{\mathrm{F}})V=(\lambda-\mu^{*})/(1+\lambda)$, which in turn promotes $T_{\mathrm{c}}$ SI .
The observed enhancement of $\gamma$ can be attributed to the expansion of the lattice constant $c$, as illustrated in Fig. 3(e). This interpretation is further supported by our band structure calculations for LaRu3Ge2. The projected band structures for LaRu3Ge2 and LaRu3Si2 are shown in Fig. 4(a),(b). Focusing on the $\Gamma-A$ line, which is parallel to the $k_{z}$-axis, we find that substitution of Si with Ge decreases the bandwidth of flat bands at $E_{\mathrm{F}}$. This indicates that the uniaxial lattice expansion along the $c$-axis by Ge doping suppresses interlayer coupling between kagome layers. Furthermore, the partial density of states in Fig. 4(c),(d) reveals a preference for the planar $4d_{xy}$ orbital over $4d_{z^{2}}$, with the latter exhibiting stronger hybridization along the $z$-direction. We note that the partial density of states (PDOS) for Ru-$4d_{xz}$ around $E_{\mathrm{F}}$ is hardly affected by Ge substitution SI . Importantly, Ge substitution leaves the crucial kagome lattice plane of Ru atoms mostly undisturbed, providing a powerful route for tuning superconductivity in LaRu3Si2. We stress the general applicability of such careful alloying studies to binary and ternary kagome metals Chen et al. (2024), with wide tunability of electronic properties, charge ordering, and superconductivity.
Acknowledgements We thank T.-h. Arima for enlightening discussions. This work was supported by JSPS KAKENHI Grants Nos. JP21K13877, JP22H04463, JP22K20348, JP23H05431, 23K13057, 24H01607, and 24H01604, as well as JST CREST Grant Nos. JPMJCR1874 and JPMJCR20T1 (Japan) and JST FOREST Grant No. JPMJFR2238 (Japan). This work was supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2426. Work at Princeton was supported by NSF, grant OAC-2118310, and by NSF through the Princeton Center for Complex Materials, a Materials Research Science and Engineering Center DMR-2011750.
Supplementary Information
S I Density functional calculation
First-principles calculations are carried out using the WIEN2k software package, which employs the full-potential linearized augmented plane wave (FP-LAPW) method with local orbitals Blaha et al. (2020). The exchange-correlation energy is treated within the Generalized Gradient Approximation (GGA) using the functional formulated by Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996) in the $P6/mmm$ structure. We employ a uniform grid of 4096 $k$-points to ensure accurate sampling of the Brillouin zone and relax the lattice constants $a$ and $c$. This yields $a=5.8260\mathrm{\AA},\ c=3.5666\mathrm{\AA}$ for LaRu3Si2 and $a=5.7967\mathrm{\AA},\ c=3.7538\mathrm{\AA}$ for hypothetical LaRu3Ge2. We calculate the band structures and the density of states in the presence of spin-orbit coupling. For the density of states (DOS) and partial density of states (PDOS), we divide the calculated values by $2$ to express them in units of $\mathrm{eV}^{-1}$ per unit cell (u.c.) per spin.
Fig. S1 shows our PDOS for the Ru-$4d_{z^{2}}$, Ru-$4d_{xy}$, and Ru-$4d_{xz}$ states of LaRu3Si2 and LaRu3Ge2. The symmetrically equivalent Ru-$4d_{x^{2}-y^{2}}$ and Ru-$4d_{yz}$ states are not shown separately. We provide projected band structures in the full Brillouin zone for LaRu3Si2 and LaRu3Ge2 in Fig. S2.
Figure S1: (color online). Partial density of states (PDOS) for Ru-$4d_{z^{2}}$, Ru-$4d_{xy}$, and Ru-$4d_{xz}$ states of LaRu3Si2 and LaRu3Ge2.
Figure S2: (color online). Projected band structure of LaRu3Si2 (a) and LaRu3Ge2 (b) in the full Brillouin zone.
S II Crystal growth
Crystals of LaRu3(Si1-xGex)2 (LRSG) are prepared by arc melting in argon atmosphere after careful evacuation to $10^{-3}\,$a. The ingot is turned and remelted at least three times, and weight loss is typically around $0.5\,\%$. To stabilize the LaRu3Si2 structure over the competing phase LaRu2Si2, excess Ru metal is added to the starting composition Barz (1980); Vandenberg and Barz (1980); Kishimoto et al. (2002); Li et al. (2011); Mielke et al. (2021). The solution limit of Ge is found to be $x=0.07$, and the required amount of Ru increases when approaching this value: We use a melt of starting composition LaRu4.5(Si1-xGex)2 for $x=0,0.02,0.05,$ and LaRu5(Si1-xGex)2 for $x=0.07$.
S III Powder X-ray diffraction
We perform powder X-ray diffraction (XRD) to confirm the formation of LaRu3(Si1-xGex)2 as well as the absence of the competing LaRu2Si2 (122). The data include reflections from nonmagnetic, elemental Ru, because Ru was added to the melt to destabilize the chemically stable 122 phase. In the main text, we neglect minor effects of subtle orthorhombic distortions in this compound. This approximation is justified by the Rietveld analysis, as the major reflections are well accounted for by space group $P6/mmm$.
S IV Magnetic susceptibility measurement
We perform magnetization measurements to detect the superconducting Meissner effect in a low magnetic field of $\mu_{0}H=0.05\,$T. Samples of mass $\sim 100\,$mg were mounted in a Quantum Design Magnetic Properties Measurement System (MPMS). We correct for the demagnetization effect in $\chi$ by approximating our roughly shaped crystals as a sphere, which has a demagnetization factor $N=1/3$. The intrinsic susceptibility is then obtained using the relation: $\chi=\chi_{\mathrm{exp}}/(1-N\chi_{\mathrm{exp}})$. Slightly smaller values than the ideal $\chi=-1$ (perfect Meissner shielding) may arise due to deviations from the spherical approximation.
S V Specific heat measurement
We measure specific heat by the relaxation technique in a Quantum Design Physical Properties Measurement System (PPMS). For each measurement, the addenda (with Apiezon N grease mounted on the stage) are measured before setting the sample for the main measurement. To eliminate contributions from non-magnetic impurity Ru metal, we subtract $\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/molK^{2}}$ and $\beta_{\mathrm{Ru}}=0.068\,\mathrm{mJ/molK^{4}}$ from the total $\gamma$ and $\beta$ values Li et al. (2011); Reese and Johnson (1970). For each value, we multiply the ratio of Ru impurity $y/3$, where $y$ represents the ratio of excess Ru, namely $y=1.5$ for $x=0,0.02,0.05$ and $y=2$ for $x=0.07$. We summarize $\gamma$ and $\beta$ values before and after the subtraction of the Ru contribution in Table 1. Errors for the fitting parameters $\beta$ and $\gamma$ in Fig. 3(d) are estimated by the bootstrap method. We estimate the density of states (DOS) at the Fermi energy, $N(E_{\mathrm{F}})$, using the relation
$N(E_{\mathrm{F}})=\frac{3\gamma}{\pi^{2}k_{\mathrm{B}}^{2}},$ (S1)
which has units of $\mathrm{J}^{-1}$. We convert this to $\mathrm{eV}^{-1}$ per unit cell (u.c.). Furthermore, since $\gamma$ is calculated assuming a spinless electronic structure, we divide by 2 to obtain $N(E_{\mathrm{F}})$ in units of $\mathrm{eV}^{-1}$ per unit cell per spin Tinkham (2004); Kriener et al. (2022).
For the superconducting state, the specific heat can be fitted by an analytical expression within the framework of the BCS theory. Here we approximate the temperature dependence to be linear below $T_{\mathrm{c}}$, which is valid for our temperature range ($>2\,$K) Gui and Cava (2022). This method allows us to precisely determine $T_{\mathrm{c}}$ by the equal-area construction, which is crucial for tracking the evolution of $T_{\mathrm{c}}$ in our crystals: We first numerically calculate the area under the curve of $\Delta C/T=C/T-C_{\mathrm{N}}/T$ below 888\,8K, denoted as $S$ [Fig. S3(a)]. We then fit $\Delta C/T$ below 666\,6K by a linear function $\Delta C/T=aT+b$. Imposing entropy conservation, the area under the curve is thus given by
$S=\frac{1}{2}(T_{\mathrm{c}}-T_{0})(aT_{\mathrm{c}}+b)-\frac{1}{2}(T_{0}-T_{\mathrm{min}})|aT_{\mathrm{min}}+b|,$ (S2)
where $T_{0}=-b/a$ is the intercept of the linear fit, and $T_{\mathrm{min}}$ is the lowest temperature at which data were taken. The first term corresponds to the area for $T_{0}\leq T\leq T_{\mathrm{c}}$ and the second to that for $T_{\mathrm{min}}\leq T\leq T_{\mathrm{c}}$, as illustrated in Fig. S3(b). Solving this equation, we analytically determine the transition temperature as
$T_{\mathrm{c}}=\frac{-b+\sqrt{b^{2}+2aS+a^{2}T_{\mathrm{min}}^{2}+2abT_{\mathrm{min}}}}{a}.$ (S3)
Furthermore, we obtain the specific heat jump at $T_{\mathrm{c}}$ as
$\Delta C(T_{\mathrm{c}})/T_{\mathrm{c}}=aT_{\mathrm{c}}+b.$ (S4)
We provide the calculated ratio $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}$ in Table 1. A slight deviation from the BCS value of $1.43$ may suggest a moderate superconducting coupling strength. Given that the Sommerfeld coefficient of pure Ru is small ($\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/mol\,K^{2}}$), the residual specific heat contribution from the Ru impurity is expected to be negligible.
To explore the effects of $\beta$ and $\gamma$ on $T_{\mathrm{c}}$, we consider the McMillan equation
$T_{\mathrm{c}}=\frac{\Theta_{\mathrm{D}}}{1.45}\exp\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right),$ (S5)
where $\Theta_{\mathrm{D}}$ is the Debye temperature, $\lambda$ is the electron-phonon coupling constant, and $\mu^{*}$ is the Coulomb pseudopotential Carbotte (1990). The electron-phonon coupling $\lambda$ is expressed in terms of the density of states at the Fermi level, $N(E_{\mathrm{F}})$, as
$N(E_{\mathrm{F}})V$ $=\frac{\lambda-\mu^{*}}{1+\lambda}$ (S6)
$=1-\frac{1+\mu^{*}}{1+\lambda},$ (S7)
where $V$ denotes the pairing potential. This relation indicates that an increase in $\gamma$ will also enhance $\lambda$, thereby promoting $T_{\mathrm{c}}$; this is assuming a relatively static value of $V$. Meanwhile, the Debye temperature $\Theta_{\mathrm{D}}$ is related to $\beta$ viz.
$\Theta_{\mathrm{D}}=\left(\frac{12\pi^{4}Nk_{B}}{5\beta}\right)^{\frac{1}{3}},$ (S8)
where $N$ is the number of atoms per formula unit. Since $\Theta_{\mathrm{D}}$ scales inversely with $\beta^{1/3}$, an increase in $\beta$ suppresses $\Theta_{\mathrm{D}}$, leading to a reduction in $T_{\mathrm{c}}$. Overall, while both $\beta$ and $\gamma$ influence $T_{\mathrm{c}}$, the observed enhancement of $T_{\mathrm{c}}$ suggests that the effect of $\gamma$ dominates. Setting $\mu^{*}=0.13$, we estimate $\lambda$ to be $0.63\sim 0.66$ from Eq. (S5), as summarized in Table 1. The enhanced electron-phonon coupling constant $\lambda>0.5$ suggests a moderately coupled superconductor. This is consistent with $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}>1.43$.
With $\lambda$ determined, we compute the bare density of states, $N_{\mathrm{bare}}(E_{\mathrm{F}})$, using
$N_{\mathrm{bare}}(E_{\mathrm{F}})=\frac{N(E_{\mathrm{F}})}{1+\lambda},$ (S9)
[…]
Figure S3: (color online). Fitting of superconducting specific heat $\Delta C=C-C_{\mathrm{N}}$ and equal-area construction. As a representative example, we show the procedure for LaRu3Si2. (a) The area under the curve (green) is calculated numerically. (b) $\Delta C/T$ is fitted by a linear approximation of the BCS expression. $T_{\mathrm{c}}$ is identified such that the gray areas equal the green area in panel (a). $T_{\mathrm{0}}$ denotes the intercept of the fit.
Table 1: Properties of LaRu3(Si1-xGex)2 probed by specific heat measurements. For the electron and phonon contributions $\gamma$ and $\beta$, values before and after subtracting a subtle contribution from Ru impurity in LaRu3(Si1-xGex)2 + $y$Ru are given. $\Delta C(T_{\mathrm{c}})$ denotes the specific heat jump at $T_{\mathrm{c}}$, and $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}=1.43$ holds for weakly coupled superconductors. $N(E_{\mathrm{F}})$ and $\lambda$ denote the DOS and the electron-phonon coupling constant, respectively.
Ge content $x$ 0 0.02 0.05 0.07
$\gamma\,\,\ \mathrm{(mJ/molK^{2})}$ 38.3 37.4 40.7 43.8
$\gamma-y/3\cdot\gamma_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{2})}$ 36.9 36.0 39.3 42.0
$\beta\,\,\mathrm{(mJ/molK^{4})}$ 0.21 0.19 0.25 0.25
$\beta-y/3\cdot\beta_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{4})}$ 0.18 0.15 0.21 0.20
$\Delta C(T_{\mathrm{c}})/\gamma T_{c}$ 1.42 1.64 1.54 1.57
$N(E_{\mathrm{F}})\,\,\mathrm{(1/eV/u.c/spin)}$ 7.82 7.64 8.33 8.90
$\lambda$ 0.64 0.63 0.66 0.66
$N(E_{\mathrm{F}})/(1+\lambda)\,\,\mathrm{(1/eV/u.c/spin)}$ 4.78 4.68 5.03 5.36